An aircraft flew 7 hours with the wind. The return trip took 8 hours with the wind. If the speed of the plane in still air is 630 miles per hour more than the speed of the wind, find the wind speed and the speed of the plane in still air.
Answer by graphmatics(170)

Let x be the speed of the wind. We note that the 7 hour flight is made with the
wind pushing the plane. So for the 7 hour flight the distance is
7*(630+x). For the 8 hour flight we have that the distance is 8*(630-x). So
7*(630+x)= 8*(630-x)
4410 + 7*x = 5040 - 8*x
15*x = 630
x = 42